### v2

# Anomalous diffusion in convergence to effective ergodicity

Abstract
Power-law exponents in the convergence to effective ergodicity is quantified for Ising-Lenz model in one dimension. Modified Thirumalai-Mountain (TM) metric for magnetisation is computed for the range of temperature values under strongly correlated dynamics. In producing evolution of TM metric over time, time-averaged dynamics is generated by using Metropolis and Glauber single-spin-flip dynamics, and ensemble-averaged dynamics with an exact solution. Superdiffusive behaviour is numerically identified in the parameter regimes studied, i.e., power-law exponents, $\alpha>1.0$.
keywords:
ergodicity, power-laws, ising models, lattice dynamics, monte carlo
articletype: Paper
1 Introduction
Dynamics of cooperation among assembly of independent units has been studied in many different context [1], such as model of magnetic material [2, 3, 4] or the state of a neuron [5]. Convergence to ergodicity in this kind of cooperation dynamics is explored recently [6] using Thirumalai-Mountain (TM) fluctuation metric [7, 8] that is adapted for Ising model’s total magnetization. In this study, we investigate the power-laws emerged from the time evolution of the rate of effective ergodic convergence under the parameters that give rise to strong correlations and anamolous diffusion in covergence. Understanding ergodicity for this circumstances is not only interesting due to fundamental importance in statistical mechanics [9] and anomalous diffusion in general [10], also for its implications in real world applications, such as understanding disruption in neural networks for dementia [11], realisation of associative memory in a solid-state device [12], nature of economic utility [13] and optical lattice dynamics [14].
The Ising model is introduced in Sec. 2. In Sec. 3, the concept of effective ergodicity will be shortly explained. Measuring convergence to ergodicity is shown in Sec. 4. Computation of power-law will be shown in Sec. 5. Summary is given in Sec. 6.
2 Ising-Lenz System
A lattice with $N$ sites, labelled as $\{s_{i}\}_{i=1}^{N}$, which can take up two values, such as $\{1,-1\}$. These values implies spin up or down in the Ising Model [2, 3, 4] or an activation in neuronal systems [5].
The total energy, Hamiltonian of the system can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
$\mathcal{H}(\{s_{i}\}_{i=1}^{N},J,H)$ $=$ $J\Big((\sum_{i=1}^{N-1}s_{i}s_{i+1})+(s_{1}s_{N})\Big)$ (1)
$+H\sum_{1}^{N}s_{i}.$
the term $s_{1}s_{N}$ is imposed by the periodic boundary conditions which provides translational invariance, making the model a closed chain of interacting units. Thermal scale is expressed with $\beta=\frac{1}{k_{B}T}$ and the corresponding coefficients for NN and external field is scaled by this,
$K=\beta J,\qquad h=\beta H.$ (2)
Analytic expression for finite size magnetisation $M_{E}(N,\beta,H)$ is obtained by using Transfer Matrix method. Whereas, time evolution of $M_{E}(N,\beta,H)$ is computed via Monte Carlo procedure with Metropolis and Glauber single-spin-flip dynamics. Where, more details can be found in the previous work [6].
3 Ergodic Dynamics
A form of ergodic dynamics suggested by Boltzmann that trajectories of many-body system will reach to phase-space partitions where its likely to be in a thermodynamic equilibrium [9]. At this point, ensemble averages and time averages of the system produce close to identical measures. This implies, for a given observable $g$ over a fixed phase-space points, ensemble averaged value can be recovered by time average values, $g(t)$ from $t_{0}$ to $t_{N}$,
$\langle g\rangle=lim_{t_{N}\to\infty}\int_{t_{0}}^{t_{N}}g(t)dt,$ (3)
where $\langle\rangle$ indicates ensemble averaged value. This kind of basic definition is not standard in the literature [15, 7, 16]. Other forms of ergodicity demands that system should visit all available phase-space partitions, which might not be possible for a finite physical system, and moreover feasibility of this type of ergodicity is questioned recently [17]. In practical terms, since partitions of the phase-space is clustered around few regions, effective ergodicity can be attained quickly [7].
On the other hand, ergodicity for the single-spin-flip dynamics, essentially a Markov process, defined as accessibility of any given state point to an other state point over time [16, 18]. In this sense, Monte Carlo procedure used in this work is ergodic by construction for long enough times.
4 Measuring Convergence
Figure 1: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
Figure 2: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics at thermodynamic equilibrium [7]. This is realised by Thirumalai-Mountain (TM) $G$-fluctuating metric [7, 8], at a given time $t_{k}$
$\Omega_{G}(t_{k})=\frac{1}{N}\sum_{j=1}^{N}\big[g_{j}(t_{k})-\langle g(t_{k})\rangle\big]^{2},$ (4)
where time average per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble average over all units,
$g_{j}(t_{k})$ $=$ $\frac{1}{k}\sum_{i=0}^{k}g_{j}(t_{i}),$ (5)
$\langle g(t_{k})\rangle$ $=$ $\frac{1}{N}\sum_{j=1}^{N}g_{j}(t_{k}).$ (6)
Consequently, the rate of ergodic convergence is measured as
$\Omega_{G}^{{}^{\prime}}=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}}$ (7)
where time average per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble average over all units,
$g_{j}(t_{k})$ $=$ $\frac{1}{k}\sum_{i=0}^{k}g_{j}(t_{i}),$ (8)
$\langle g(t_{k})\rangle$ $=$ $\frac{1}{N}\sum_{j=1}^{N}g_{j}(t_{k}).$ (9)
Consequently, the rate of ergodic convergence is measured as
$\Omega_{G}^{{}^{\prime}}=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}}$ (10)
where diffusion coeffient expressed as $D_{G}$. Similar measure of ergodicity is used in simple liquids [7, 19], and earthquake fault networks [20, 21].
Using $\Omega_{G}$ for total magnetization in Ising Model at time $t_{k}$ as a function of temperature and external field values reads
$\Omega_{M}(t_{k},N,\beta,h)$ $=$ $\big[M_{T}(t_{k})-M_{E}\big]^{2},$ (11)
$M_{T}$ $=$ $\frac{1}{k}\sum_{i=0}^{k}M(t_{i}),$
where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time and ensemble averaged total magnetization. Note that exact expression for $M_{E}$ is used [6]. Hence in computing $\Omega_{M}(t_{k},N,\beta,h)$, we generate Metropolis and Glauber single-spin-flip dynamics for $N=512$ cooperating units, a range of Boltzmann factor values $\beta=[0.5,1.5]$ with a spacing of $0.025$ and non-zero external field values $H=1.0$ with setting short-range interaction strength to $J=1.0$ [22, 23]. Time evolution of inverse rate, $\Omega_{G}^{-1}$ is given in Figure 1a and Figure 1b for $N=512$. After an initial period, inverse rate value decreases rapidly. At higher temperatures, inverse rate behaves differently, we quantify this using power-laws.
Figure 3: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively. Sharp increase at around $\beta=1.4$.
Figure 4: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively. Sharp increase at around $\beta=1.4$.
5 Power laws
Inverse rate $\Omega_{G}^{{}^{\prime}-1}$ in the effective ergodic measure scales with Monte Carlo time in a power-law fashion,
$\Omega_{G}^{{}^{\prime}-1}\sim t^{-\alpha}$
where $\alpha$ is the the exponent and $t$ is the Monte Carlo time. While, power laws in complex systems and computing them on a given emprical data are studied in depth [24, 25]. We compute power-law exponents of evolution of inverse rate over Monte Carlo time for the parameters given at Figure LABEL:fig:metropolis and LABEL:fig:glauber. Identification where to start to compute $\alpha$ is achived by Kolmogorov-Smirnov test [26]. Temperature dependence of $\alpha$ is reported in Figure 3a and Figure 3b for $N=512$, where $\alpha>1.0$ is observed leading to superdiffusive behaviour. We also observe that there is a sharp increase for $\alpha$ at around $\beta=1.4$, along with a stable value in the range $\beta=(0.8,1.4)$.
6 Summary
Understanding how ergodicity is reached emprically plays a vital role both theoretically and from practical standpoint. We have identified superdiffusive behaviour and power-law exponents for the rate of ergodic convergence in Ising Model under parameters give rise to strong correlation, using the modified Thirumalai-Mountain (TM) metric for the total magnetization. We conclude that combination of stronger temperature values give rise to lower exponents, attributed to high external field.

### v3

@@
 Abstract
-Power-law exponents in the convergence to effective ergodicity is quantified for Ising-Lenz model in one dimension.
-Modified Thirumalai-Mountain (TM) metric for magnetisation is computed for the range of temperature values under strongly correlated dynamics.
-In producing evolution of TM metric over time, time-averaged dynamics is generated by using Metropolis and Glauber single-spin-flip dynamics, and ensemble-averaged dynamics with an exact solution.
-Superdiffusive behaviour is numerically identified in the parameter regimes studied, i.e., power-law exponents, $\alpha>1.0$.
+The nature of diffusion is usually studied for particles or dynamics generating trajectories over time.
+Similar in principle, these studies can be executed in tracking how a given function of observable properties evolve over time akin to a system’s particle motion, so-called functional-diffusion.
+This is not the same as systems’ own trajectories but can be considered as a meta-trajectory.
+Following this idea, we measure how an approach to ergodicity evolves over time for the observed magnetization of a full Ising model with an external field.
+We compute the functional’s diffusive behavior depending on a range of temperatures via Metropolis and Glauber single-spin-flip dynamics.
+System’s ensemble-averaged dynamics are computed using expressions from the exact solution.
+Power-laws on the approach to ergodicity provide the classification of anomalies in the functional-diffusion, demonstrating non-linear super-diffusive behavior over different temperature ranges.
 keywords:
-ergodicity, power-laws, ising models, lattice dynamics, monte carlo
+ergodicity, power-laws, Ising models, lattice dynamics, monte carlo, functional-diffusion
 articletype: Paper
 1 Introduction
-Dynamics of cooperation among assembly of independent units has been studied in many different context [1], such as model of magnetic material [2, 3, 4] or the state of a neuron [5].
-Convergence to ergodicity in this kind of cooperation dynamics is explored recently [6] using Thirumalai-Mountain (TM) fluctuation metric [7, 8] that is adapted for Ising model’s total magnetization.
-In this study, we investigate the power-laws emerged from the time evolution of the rate of effective ergodic convergence under the parameters that give rise to strong correlations and anamolous diffusion in covergence.
-Understanding ergodicity for this circumstances is not only interesting due to fundamental importance in statistical mechanics [9] and anomalous diffusion in general [10], also for its implications in real world applications, such as understanding disruption in neural networks for dementia [11], realisation of associative memory in a solid-state device [12], nature of economic utility [13] and optical lattice dynamics [14].
-The Ising model is introduced in Sec.
+Brownian motion is probably one of the landmark concepts in statistical physics attracted Einstein’s interest early on [1].
+Its importance in formulating statistical mechanics recently reviewed within its 250th anniverssary publications [2].
+Major observable in tracing Brownian motion appears as how accumulated displacement curve behaves over time, i.e., linear relationship without intercept corresponding to normal diffusion.
+Along these lines, dynamics of cooperation among assembly of independent units has been studied in this context [3], such as model of magnetic material [4, 5, 6] and the state of a neuron [7, 8].
+Measuring the meta-diffusion in this discrete cases are our core premise in an interesting setting: If the displacement curve of a functional of a function of trajectories shows a power-law scaling over time other than the exponent $1.0$, then the behaviour is called anomalous diffusion [9].
+This setting not only yields to mathematically challenging consequences, provides insights into characteristics of the Ising model.
+Important to note that, described scenario does not apply to trajectories, rather a functional of trajectories.
+This distinction is the core driver in our work that a functional on the function that explains convergence to ergodicity for total magnetisation is the diffusing, rather than the discrete units of the Ising model.
+Convergence to ergodicity in this kind of cooperation dynamics is explored and established [10] using Thirumalai-Mountain (TM) fluctuation metric [11, 12] that is adapted for the Ising model’s total magnetization.
+In this study, we investigate the power-laws that emerged from the time evolution of the rate of effective ergodic convergence under the parameters that give rise to strong correlations and anomalous diffusion in convergence.
+Understanding ergodicity for these circumstances is not only interesting due to fundamental importance in statistical mechanics [13] and anomalous diffusion in general but also for its implications in real-world applications, such as understanding disruption in neural networks for dementia [14], realization of associative memory in a solid-state device [15], nature of economic utility [16] and optical lattice dynamics [17].
+The formulation of the Ising model is reviewed in Sec.
 2.
 In Sec.
-3, the concept of effective ergodicity will be shortly explained.
+3, the concept of effective ergodicity is explained.
 Measuring convergence to ergodicity is shown in Sec.
-4.
+4; this corresponds to the functional-diffusion concept we introduced.
 Computation of power-law will be shown in Sec.
 5.
-Summary is given in Sec.
+The conclusion is given in Sec.
 6.
 2 Ising-Lenz System
-A lattice with $N$ sites, labelled as $\{s_{i}\}_{i=1}^{N}$, which can take up two values, such as $\{1,-1\}$.
-These values implies spin up or down in the Ising Model [2, 3, 4] or an activation in neuronal systems [5].
-The total energy, Hamiltonian of the system can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
-$\mathcal{H}(\{s_{i}\}_{i=1}^{N},J,H)$ $=$ $J\Big((\sum_{i=1}^{N-1}s_{i}s_{i+1})+(s_{1}s_{N})\Big)$ (1)
-$+H\sum_{1}^{N}s_{i}.$
-the term $s_{1}s_{N}$ is imposed by the periodic boundary conditions which provides translational invariance, making the model a closed chain of interacting units.
-Thermal scale is expressed with $\beta=\frac{1}{k_{B}T}$ and the corresponding coefficients for NN and external field is scaled by this,
+A lattice with $N$ sites, labelled as $\{s_{i}\}_{i=1}^{N}$ which can take two values, such as $\{1,-1\}$.
+These values imply spin up or down in the Ising Model [4, 5, 6] or an activation in neuronal systems [7, 8].
+The total energy, the Hamiltonian of the system, can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
+$\mathcal{H}({s_{i}}{i=1}^{N},J,H)=J\Big((\sum_{i=1}^{N-1}s_{i}s_{i+1})+(s_{1}s_{N})\Big)+H\sum_{1}^{N}s_{i}.$ (1)
+The term $s_{1}s_{N}$ is imposed by the periodic boundary conditions, which provide translational invariance, making the model a closed chain of interacting units.
+The thermal scale is expressed with $\beta=\frac{1}{k_{B}T}$ and the corresponding coefficients for NN and external field are scaled by this,
 $K=\beta J,\qquad h=\beta H.$ (2)
-Analytic expression for finite size magnetisation $M_{E}(N,\beta,H)$ is obtained by using Transfer Matrix method.
-Whereas, time evolution of $M_{E}(N,\beta,H)$ is computed via Monte Carlo procedure with Metropolis and Glauber single-spin-flip dynamics.
-Where, more details can be found in the previous work [6].
+The analytic expression for finite size magnetisation $M_{E}(N,\beta,H)$ is obtained by using the Transfer Matrix method.
+Whereas the time evolution of $M_{E}(N,\beta,H)$ i is computed via the Monte Carlo procedure with Metropolis and Glauber single-spin-flip dynamics.
+More details can be found in the previous work [10].
+Figure 1: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
 3 Ergodic Dynamics
-A form of ergodic dynamics suggested by Boltzmann that trajectories of many-body system will reach to phase-space partitions where its likely to be in a thermodynamic equilibrium [9].
+A form of ergodic dynamics suggested by Boltzmann that trajectories of a many-body system will reach phase-space partitions where it’s likely to be in a thermodynamic equilibrium [13].
 At this point, ensemble averages and time averages of the system produce close to identical measures.
-This implies, for a given observable $g$ over a fixed phase-space points, ensemble averaged value can be recovered by time average values, $g(t)$ from $t_{0}$ to $t_{N}$,
+This implies for a given observable $g$ over a fixed phase-space point, ensemble-averaged value can be recovered by time-averaged values, $g(t)$ from $t_{0}$ to $t_{N}$,
 $\langle g\rangle=lim_{t_{N}\to\infty}\int_{t_{0}}^{t_{N}}g(t)dt,$ (3)
-where $\langle\rangle$ indicates ensemble averaged value.
+where $\langle\rangle$ indicates ensemble-averaged value.
 This kind of basic definition is not standard in the literature [18, 11, 19].
-Other forms of ergodicity demands that system should visit all available phase-space partitions, which might not be possible for a finite physical system, and moreover feasibility of this type of ergodicity is questioned recently [17].
-In practical terms, since partitions of the phase-space is clustered around few regions, effective ergodicity can be attained quickly [7].
+Other forms of ergodicity demand that system should visit all available phase-space partitions, which might not be possible for a finite physical system, and moreover feasibility of this type of ergodicity is questioned [20].
+In practical terms, since partitions of the phase-space are clustered around few regions, effective ergodicity can be attained quickly [11].
 On the other hand, ergodicity for the single-spin-flip dynamics, essentially a Markov process, defined as accessibility of any given state point to another state point over time [19, 21].
 In this sense, Monte Carlo procedure used in this work is ergodic by construction for long enough times.
-4 Measuring Convergence
-Figure 1: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
+4 Ergodicity convergence as functional-diffusion
 Figure 2: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
-Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics at thermodynamic equilibrium [7].
-This is realised by Thirumalai-Mountain (TM) $G$-fluctuating metric [7, 8], at a given time $t_{k}$
+Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics in thermodynamic equilibrium [11].
+This is realised by the Thirumalai-Mountain (TM) $G$-fluctuating metric [11, 12], at a given time $t_{k}$
 $\Omega_{G}(t_{k})=\frac{1}{N}\sum_{j=1}^{N}\big[g_{j}(t_{k})-\langle g(t_{k})\rangle\big]^{2},$ (4)
-where time average per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble average over all units,
+Functional-diffusion occurs for the development of a function of an observable over time, $F[O(t)]$, measured as its displacement from the initial conditions.In this sense, time-evolution of inverse of $\Omega_{G}(t_{k})$ can be considered as it is in functional-diffusion
+Figure 3: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
+Sharp increase at around $\beta=1.4$.
+where time-averaged per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble-averaged over all units,
 $g_{j}(t_{k})$ $=$ $\frac{1}{k}\sum_{i=0}^{k}g_{j}(t_{i}),$ (5)
@@
 $\Omega_{G}^{{}^{\prime}}=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}}$ (7)
-where time average per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble average over all units,
-$g_{j}(t_{k})$ $=$ $\frac{1}{k}\sum_{i=0}^{k}g_{j}(t_{i}),$ (8)
-$\langle g(t_{k})\rangle$ $=$ $\frac{1}{N}\sum_{j=1}^{N}g_{j}(t_{k}).$ (9)
-Consequently, the rate of ergodic convergence is measured as
-$\Omega_{G}^{{}^{\prime}}=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}}$ (10)
-where diffusion coeffient expressed as $D_{G}$. Similar measure of ergodicity is used in simple liquids [7, 19], and earthquake fault networks [20, 21].
-Using $\Omega_{G}$ for total magnetization in Ising Model at time $t_{k}$ as a function of temperature and external field values reads
+where diffusion coefficient expressed as $D_{G}$. A similar measure of ergodicity is used in simple liquids [11, 22], and earthquake fault networks [23, 24].
+Using $\Omega_{G}$ for the total magnetization in Ising Model at time $t_{k}$ as a function of temperature and external field values is expressed as
 $\Omega_{M}(t_{k},N,\beta,h)$ $=$ $\big[M_{T}(t_{k})-M_{E}\big]^{2},$ (8)
 $M_{T}$ $=$ $\frac{1}{k}\sum_{i=0}^{k}M(t_{i}),$
-where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time and ensemble averaged total magnetization.
+where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time-averaged and ensemble-averaged total magnetization.
 Note that exact expression for $M_{E}$ is used [10].
-Hence in computing $\Omega_{M}(t_{k},N,\beta,h)$, we generate Metropolis and Glauber single-spin-flip dynamics for $N=512$ cooperating units, a range of Boltzmann factor values $\beta=[0.5,1.5]$ with a spacing of $0.025$ and non-zero external field values $H=1.0$ with setting short-range interaction strength to $J=1.0$ [22, 23].
-Time evolution of inverse rate, $\Omega_{G}^{-1}$ is given in Figure 1a and Figure 1b for $N=512$.
+Hence, in computing $\Omega_{M}(t_{k},N,\beta,h)$, we generate Metropolis and Glauber single-spin-flip dynamics for $N=512$ cooperating units, a range of Boltzmann factor values $\beta=[0.5,1.5]$ with a spacing of $0.025$ and non-zero external field values $H=1.0$ with setting short-range interaction strength to $J=1.0$ [25, 26].
+The time evolution of inverse rate, $\Omega_{G}^{-1}$, is given in Figure 1a and Figure 1b for $N=512$.
 After an initial period, inverse rate value decreases rapidly.
-At higher temperatures, inverse rate behaves differently, we quantify this using power-laws.
-Figure 3: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
-Sharp increase at around $\beta=1.4$.
+At higher temperatures, inverse rate behaves differently.
+Similar results are produced for $N=1024$ on Figure 2a and Figure 2b.
 Figure 4: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
 Sharp increase at around $\beta=1.4$.
-5 Power laws
+5 Anomaly in the functional-diffusion
 Inverse rate $\Omega_{G}^{{}^{\prime}-1}$ in the effective ergodic measure scales with Monte Carlo time in a power-law fashion,
 $\Omega_{G}^{{}^{\prime}-1}\sim t^{-\alpha}$
-where $\alpha$ is the the exponent and $t$ is the Monte Carlo time.
-While, power laws in complex systems and computing them on a given emprical data are studied in depth [24, 25].
-We compute power-law exponents of evolution of inverse rate over Monte Carlo time for the parameters given at Figure LABEL:fig:metropolis and LABEL:fig:glauber.
+where $\alpha$ is the exponent and $t$ is the Monte Carlo time.
+While power laws in complex systems and computing them on a given empirical data are studied in depth [27, 28].
+We compute power-law exponents of evolution of inverse rate over Monte Carlo time.
 Identification where to start to compute $\alpha$ is achived by Kolmogorov-Smirnov test [29].
@@
 We also observe that there is a sharp increase for $\alpha$ at around $\beta=1.4$, along with a stable value in the range $\beta=(0.8,1.4)$.
-6 Summary
-Understanding how ergodicity is reached emprically plays a vital role both theoretically and from practical standpoint.
-We have identified superdiffusive behaviour and power-law exponents for the rate of ergodic convergence in Ising Model under parameters give rise to strong correlation, using the modified Thirumalai-Mountain (TM) metric for the total magnetization.
-We conclude that combination of stronger temperature values give rise to lower exponents, attributed to high external field.
+Due to different relaxation regimes, results deviate for $N=1024$, as seen in Figure 4a and Figure 4b, a parabola-like behaviour for superdussion strengths is observed over a temperature dependence.
+6 Conclusion
+A new concept of functional-diffusion is introduced via a canonical example on Ising model i.e., approach to ergodicity for total magnetisation.
+Superdiffusive regimes are identified based on different temperature ranges, via computation of power-laws of the inverse fluctuating metric for ergodicity.
+The results demonstrate size dependence on the relaxation dynamics and their anomalous diffusion for the functional cases.
+This concept would be a pedagogical test-bed for the extentions of diffusive behaviour beyond particle-trajectories.
+Supplement
+We provide two notebooks for reproducing the results, data_analysis.ipynb and data_generate.ipynb. Dataset is available on Zenodo [30]

### v4

@@
 System’s ensemble-averaged dynamics are computed using expressions from the exact solution.
-Power-laws on the approach to ergodicity provide the classification of anomalies in the functional-diffusion, demonstrating non-linear super-diffusive behavior over different temperature ranges.
+Power-laws on the approach to ergodicity provide the classification of anomalies in the functional-diffusion, demonstrating non-linear anomalous behavior over different temperature and field ranges.
 keywords:
-ergodicity, power-laws, Ising models, lattice dynamics, monte carlo, functional-diffusion
+ergodicity, power-laws, Ising models, lattice dynamics, monte carlo, functional-diffusion, Glauber dynamics, Metropolis dynamics
 articletype: Paper
@@
 Brownian motion is probably one of the landmark concepts in statistical physics attracted Einstein’s interest early on [1].
-Its importance in formulating statistical mechanics recently reviewed within its 250th anniverssary publications [2].
+Its importance in formulating statistical mechanics recently reviewed within its 250th anniversary publications [2].
 Major observable in tracing Brownian motion appears as how accumulated displacement curve behaves over time, i.e., linear relationship without intercept corresponding to normal diffusion.
-Along these lines, dynamics of cooperation among assembly of independent units has been studied in this context [3], such as model of magnetic material [4, 5, 6] and the state of a neuron [7, 8].
-Measuring the meta-diffusion in this discrete cases are our core premise in an interesting setting: If the displacement curve of a functional of a function of trajectories shows a power-law scaling over time other than the exponent $1.0$, then the behaviour is called anomalous diffusion [9].
-This setting not only yields to mathematically challenging consequences, provides insights into characteristics of the Ising model.
-Important to note that, described scenario does not apply to trajectories, rather a functional of trajectories.
-This distinction is the core driver in our work that a functional on the function that explains convergence to ergodicity for total magnetisation is the diffusing, rather than the discrete units of the Ising model.
+If the displacement curve of trajectories shows a power-law scaling over time other than the linear relationship, then the behaviour is called anomalous diffusion [3, 4, 5].
+There is a recent surge in interest in using machine learning techniques in the analysis of anomalous diffusion data [6, 7, 8, 9], for example, the characterisation using transformer architectures is quite novel given the larger interests in attention mechanism in machine learning [10].
+Along these lines, the dynamics of cooperation among assembly of independent units has been studied in this context [11], such as model of magnetic material [12, 13, 14] and the state of a neuron [15, 16].
+Measuring the diffusion behaviour in this discrete cases would not only yields to mathematically challenging consequences, provides insights into characteristics of the Ising model.
+In our described scenario does not apply to trajectories, rather a functional of trajectories.
+This distinction is the core driver in our work that a functional on the function that explains convergence to ergodicity for total magnetisation is the diffusing, rather than the trajectory of the discrete units of the Ising model.
 Convergence to ergodicity in this kind of cooperation dynamics is explored and established [17] using Thirumalai-Mountain (TM) fluctuation metric [18, 19] that is adapted for the Ising model’s total magnetization.
@@
 Computation of power-law will be shown in Sec.
-5.
+5 and our experimental results in Sec 6.
 The conclusion is given in Sec.
-6.
+7.
 2 Ising-Lenz System
-A lattice with $N$ sites, labelled as $\{s_{i}\}_{i=1}^{N}$ which can take two values, such as $\{1,-1\}$.
+A lattice with $N$ sites, labeled as $\{s_{i}\}_{i=1}^{N}$ which can take two values, such as $\{1,-1\}$.
 These values imply spin up or down in the Ising Model [12, 13, 14] or an activation in neuronal systems [15, 16].
@@
 More details can be found in the previous work [17].
-Figure 1: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
+Figure 1: Diagnostic plots of $40$ different runs for $N=1024$ with Glauber dynamics for the evolution of $K(t)$ log-log regression (a) A field at $H=1.0$ and inverse temperature $\beta=1.16$ (b) A field at $H=1.4$ and inverse temperature $\beta=1.0$.
+We identify optimal time starting fit with minimal Kolmogorov-Simirnov statistic on a grid-search.
+Figure 2: Uncertainty quantified two power-laws’s $\alpha$ exponents over temperature ranges (a) the time-evolution of $K(t)$, via log-log regressions.
+(b) the distribution of $\Gamma(t)$ via analytical expressions from Newman et.
+al..
+Here we plot absolute value as $\alpha$.
+In the fit, functional form is negative $\alpha$.
 3 Ergodic Dynamics
@@
 4 Ergodicity convergence as functional-diffusion
-Figure 2: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
 Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics in thermodynamic equilibrium [18].
-This is realised by the Thirumalai-Mountain (TM) $G$-fluctuating metric [11, 12], at a given time $t_{k}$
+This is realized by the Thirumalai-Mountain (TM) $G$-fluctuating metric [18, 19].
+Applications of ergodicity with TM metric is quite widely used also with recent interests, biomolecular simulations [29], physical chemistry and machine learning landscapes [30], seismology [31, 32, 33, 34, 35, 36], neuromorphic computing [37, 38] and artificial spin ice [39, 40].
+TM metric at a given time $t_{k}$ reads,
 $\Omega_{G}(t_{k})=\frac{1}{N}\sum_{j=1}^{N}\big[g_{j}(t_{k})-\langle g(t_{k})\rangle\big]^{2},$ (4)
 Functional-diffusion occurs for the development of a function of an observable over time, $F[O(t)]$, measured as its displacement from the initial conditions.In this sense, time-evolution of inverse of $\Omega_{G}(t_{k})$ can be considered as it is in functional-diffusion
-Figure 3: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
-Sharp increase at around $\beta=1.4$.
 where time-averaged per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble-averaged over all units,
@@
 Consequently, the rate of ergodic convergence is measured as
-$\Omega_{G}^{{}^{\prime}}=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}}$ (7)
-where diffusion coefficient expressed as $D_{G}$. A similar measure of ergodicity is used in simple liquids [11, 22], and earthquake fault networks [23, 24].
+$\Gamma_{G}(t)=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}},$ (7)
+where diffusion coefficient expressed as $D_{G}$.
+A similar measure of ergodicity is used in simple liquids [18, 41], and earthquake fault networks [42, 43].
+The inverse rate of ergodic convergence can also be defined by $K_{G}(t)$,
+$K_{G}(t)=\frac{\Omega_{G}(0)}{\Omega_{G}(t)}\to t\cdot D_{G}.$ (8)
 Using $\Omega_{G}$ for the total magnetization in Ising Model at time $t_{k}$ as a function of temperature and external field values is expressed as
@@
 where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time-averaged and ensemble-averaged total magnetization.
+Since, ensemble averaged is computed analytically, this makes our approach using a modified TM-metric, leading to more accurate assessment of the ensemble averaging.
 Note that exact expression for $M_{E}$ is used [17].
-Hence, in computing $\Omega_{M}(t_{k},N,\beta,h)$, we generate Metropolis and Glauber single-spin-flip dynamics for $N=512$ cooperating units, a range of Boltzmann factor values $\beta=[0.5,1.5]$ with a spacing of $0.025$ and non-zero external field values $H=1.0$ with setting short-range interaction strength to $J=1.0$ [25, 26].
-The time evolution of inverse rate, $\Omega_{G}^{-1}$, is given in Figure 1a and Figure 1b for $N=512$.
-After an initial period, inverse rate value decreases rapidly.
-At higher temperatures, inverse rate behaves differently.
-Similar results are produced for $N=1024$ on Figure 2a and Figure 2b.
-Figure 4: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
-Sharp increase at around $\beta=1.4$.
-5 Anomaly in the functional-diffusion
-Inverse rate $\Omega_{G}^{{}^{\prime}-1}$ in the effective ergodic measure scales with Monte Carlo time in a power-law fashion,
-$\Omega_{G}^{{}^{\prime}-1}\sim t^{-\alpha}$
-where $\alpha$ is the exponent and $t$ is the Monte Carlo time.
-While power laws in complex systems and computing them on a given empirical data are studied in depth [27, 28].
-We compute power-law exponents of evolution of inverse rate over Monte Carlo time.
-Identification where to start to compute $\alpha$ is achived by Kolmogorov-Smirnov test [29].
-Temperature dependence of $\alpha$ is reported in Figure 3a and Figure 3b for $N=512$, where $\alpha>1.0$ is observed leading to superdiffusive behaviour.
-We also observe that there is a sharp increase for $\alpha$ at around $\beta=1.4$, along with a stable value in the range $\beta=(0.8,1.4)$.
-Due to different relaxation regimes, results deviate for $N=1024$, as seen in Figure 4a and Figure 4b, a parabola-like behaviour for superdussion strengths is observed over a temperature dependence.
+5 Functional-diffusion: Two power laws
+We would like to investigate the behaviour of the functional-diffusion in two power laws.
+The first-one is similar to the one investigated for displacement vector from the literature, but over ergodicity convergence function $K$ we defined above,
+$K_{G}(t)\to C\cdot t^{\alpha}.$ (11)
+whereby $C$ is the generalized diffusion coefficient, $t$ is the Monte Carlo time and for the positive $\alpha$ exponent.
+We call this time power-laws.
+The second type of power-law we seek is over the distribution of $\Gamma(t)$,
+$P(\Gamma_{G}(t))\to\Gamma_{G}(t)^{-\alpha}.$ (12)
+This is identified as distributional power-laws, whereby this is related to Lèvy flights or jump distributions [4], we consider cumulative jumps of approach to ergodicity here.
+We will explicitly state which power-laws we are working on, not to confuse $\alpha$-s.
+The power laws in complex systems and computing them on a given empirical data are studied in depth using techniques pioneered by Newman et.
+al.
+[44, 45].
+We did an extensive bias corrected bootstrapping for determining uncertainties on scaling exponents, diffusion coefficient, fit diagnostics, KS-distances (Kolmogorov-Simirnov Statistic) and adjusted R-squared, and autocorrelation times.
+KS-distance was critical in determining where to start to perform log-log regression depending on the parameters, this is done in an automated way with finding minimum KS-statistic over a search grid of time values.
+Visual inspection diagnostic is done for every parameters and diagnostic visualisations are developed and automated.
+We also computed autocorrelations times using time self-correlations of averaged magnetisation, via location of the plateau of the correlation times $C(t)$ of averaged-magnetisations,
+$C(t_{k})=\frac{1}{t_{k}}\sum_{i=0}^{k}M(t_{0})M(t_{i})$
+. We identify that relaxation changes in the regions where anomalous convergence appears.
+A Finite Size-Scaling (FSS) analysis could show a universality of the results depending on the number of discrete units and temperatures for the Ising model.
+We made the Following FSS assumptions, for temperature and size dependence
+$u(\beta)=(\beta-\beta_{c})N^{a}$
+scaling function $f(u)=A+Bu^{c}$, then power-law exponents $\alpha$ with the critical temperature $\beta_{c}$,
+$\alpha(T,N)=N^{b}f(u)$
+We identified the FSS exponents $a$, $b$, $c$ and coefficients $A$, $B$ with nonlinear optimization on the power-laws over temperatures for time power-laws.
+Figure 3: Uncertainty quantified two power-laws’s $\alpha$ exponents over field ranges at $\beta=1.0$ (a) the time-evolution of $K(t)$, via log-log regressions.
+(b) the distribution of $\Gamma(t)$ via analytical expressions from Newman et.
+al..
+Here we plot absolute value as $\alpha$.
+In the fit, functional form is negative $\alpha$.
+Figure 4: Uncertainty quantified power-laws of the time-evolution of $K(t)$, via log-log regressions, diffusion coefficients (a) Over temperature ranges. (b) Over field ranges.
+6 Experimental Results
+We run extensive Metropolis and Glauber dynamics for the full Ising model and measure the rate of ergodic convergence and its inverse, $\Gamma(t)$ and its inverse $K(t)$.
+Here we present the result with our findings.
+An of evolution of $K(t)$ is shown for 40 independent runs in Figure 1a and 1b for different temperature and field conditions, with it’s diagnostic of where we start fitting via optimal KS distance analysis.
+These plots served as a visualisal diagnostics in ensuring each temperature, size or field cases are inspected that the fitted log-log regression represents indeed a power-law region within the simulated MC time, whereby we record $\Gamma(t)$ are acceptance times.
+Figure 5: Uncertainty quantified autocorrelation times over $200$ runs.
+(a) A diagnostic plot for identification of autocorrelation time.
+(b) Autocorrelation times over temperature ranges and three different sizes.
+ikBT alphaMean alphaLower alphaUpper adjR2Mean adjR2Lower adjR2Upper
+$0.50$ $0.0179$ $0.0145$ $0.0212$ $0.7133$ $0.6264$ $0.7830$
+$0.56$ $0.0416$ $0.0368$ $0.0448$ $0.8870$ $0.8179$ $0.9186$
+$0.62$ $0.0872$ $0.0818$ $0.0923$ $0.9710$ $0.9617$ $0.9769$
+$0.68$ $0.1784$ $0.1726$ $0.1852$ $0.9815$ $0.9775$ $0.9847$
+$0.74$ $0.3538$ $0.3462$ $0.3608$ $0.9897$ $0.9879$ $0.9913$
+$0.80$ $0.6524$ $0.6407$ $0.6617$ $0.9939$ $0.9934$ $0.9946$
+$0.86$ $1.1030$ $1.0812$ $1.1277$ $0.9969$ $0.9964$ $0.9972$
+$0.92$ $1.4959$ $1.4892$ $1.5030$ $0.9995$ $0.9994$ $0.9995$
+$0.98$ $1.7725$ $1.7665$ $1.7784$ $0.9999$ $0.9999$ $0.9999$
+$1.04$ $1.8869$ $1.8780$ $1.8947$ $0.9998$ $0.9997$ $0.9998$
+$1.10$ $1.8857$ $1.8754$ $1.8942$ $0.9995$ $0.9994$ $0.9996$
+$1.16$ $1.8223$ $1.8110$ $1.8334$ $0.9992$ $0.9990$ $0.9993$
+$1.22$ $1.7088$ $1.6951$ $1.7257$ $0.9990$ $0.9986$ $0.9991$
+$1.28$ $1.5846$ $1.5678$ $1.6033$ $0.9983$ $0.9979$ $0.9986$
+$1.34$ $1.4132$ $1.3864$ $1.4342$ $0.9966$ $0.9942$ $0.9975$
+$1.40$ $1.1320$ $1.0944$ $1.1654$ $0.9876$ $0.9832$ $0.9907$
+Table 1: Diagnostics summary for the power-law fit for the Glauber dynamics for log-log regression.
+Power-law exponents alphaalpha for different sizes $N=512,1024,1536$ are identified for inverse temperature ranges.
+The log-log regression results are shown in Figure 2a for time power-laws, i.e., $K(t)$ observations over time.
+This is the functional-diffusion power-laws aligned with the conventional displacement power-laws.
+We observe anomalous behaviour super-diffusive and sub-diffusive regions, and normal diffusion region being in between.
+power-law has a different nature as shown in Figure 2b, it doesn’t necessarily represents diffusion behaviour, rather how Monte Carlo dynamics evolve.
+From simulations perspective this captures the dynamical behaviour due to Glauber spin-flip-dynamics, validating that anomalies present in the dynamics as well.
+More sophisticated cluster-flips maybe required for slow dynamics regions, such as the Swendsen-Wang [46] dynamics.
+Uncertainty estimates are achived using bias corrected bootstrapping for reliable confidence interval identification [47, 48].
+We demonstrate a set of uncertainty quantification for Glauber dynamics $K(t)$ power-laws log-log plots in Table 1.
+The similar observations have been made for exponents $\alpha$ the field variations in Figures 3a and 3b.
+Generalized diffusion coefficients over time validates the same observations, 4a and 4b.
+Identification of the autocorrelation times from time correlations of average magnetisation.
+This indicates the different relaxation times matching with diffusion regimes we observed, in Figures 5a and 5b.
+df type_power_law N D pvalue parameter
+1 Distribution Temperature Range $512$ $0.1875$ $0.9522$ alpha scale
+2 Distribution Temperature Range $1024$ $0.1875$ $0.9522$ alpha scale
+3 Distribution Temperature Range $1536$ $0.1250$ $0.9998$ alpha scale
+4 Distribution Field Range $512$ $0.2500$ $0.7164$ alpha scale
+5 Distribution Field Range $1024$ $0.0625$ $1.0000$ alpha scale
+6 Distribution Field Range $1536$ $0.1250$ $0.9998$ alpha scale
+7 Time Temperature Range $512$ $0.0625$ $1.0000$ alpha scale
+8 Time Temperature Range $1024$ $0.0625$ $1.0000$ alpha scale
+9 Timer Temperature Range $1536$ $0.1250$ $0.9998$ alpha scale
+10 Time Temperature Range $512$ $0.1250$ $0.9998$ C diffusion
+11 Time Temperature Range $1024$ $0.1250$ $0.9998$ C diffusion
+12 Timer Temperature Range $1536$ $0.1250$ $0.9998$ C diffusion
+13 Time Field Range $512$ $0.0625$ $1.0000$ alpha scale
+14 Time Field Range $1024$ $0.1875$ $0.9522$ alpha scale
+15 Time Field Range $1536$ $0.1250$ $0.9998$ alpha scale
+16 Time Field Range $512$ $0.0625$ $1.0000$ C diffusion
+17 Time Field Range $1024$ $0.1250$ $0.9998$ C diffusion
+18 Time Field Range $1536$ $0.1875$ $0.9522$ C diffusion
+Table 2: Diagnostic summary for comparing power-laws on different ranges, with Glauber and Metropolis dynamics via two-sided KS test.
+Time means $K(t)$ power-laws and distribution means $P(\Gamma(t))$ power-laws.
+Finite-size scaling parameters are also computed with nonlinear optimization.
+We found that $N=1024,1536$ behaves universally.
+It is also establised that results are statistically significant for both Metropolis and Glauber dynamics 2.
 7 Conclusion
-A new concept of functional-diffusion is introduced via a canonical example on Ising model i.e., approach to ergodicity for total magnetisation.
-Superdiffusive regimes are identified based on different temperature ranges, via computation of power-laws of the inverse fluctuating metric for ergodicity.
-The results demonstrate size dependence on the relaxation dynamics and their anomalous diffusion for the functional cases.
-This concept would be a pedagogical test-bed for the extentions of diffusive behaviour beyond particle-trajectories.
-Supplement
-We provide two notebooks for reproducing the results, data_analysis.ipynb and data_generate.ipynb. Dataset is available on Zenodo [30]
+A new concept of functional-diffusion is introduced via a canonical example on Ising model i.e., approach to ergodicity for total magnetization.
+Superdiffusive regimes are identified based on different temperature and field ranges, via computation of power-laws of the forward and inverse fluctuating metric for ergodicity.
+Quite comprehensive results demonstrate for the first time anomalous convergences to ergodicity quantitatively with power-laws.
+This concept would be a pedagogical test-bed for the extensions of diffusive behaviour beyond particle-trajectories, so called the functional diffusion for full Ising model and increase the awareness on possible anomalous convergence of metric functionals.
+Data Availability
+Main diagnostic tables, all data generation and analysis R notebooks are available in public domain at a Zenodo repository [49].
+Ackowlegements
+We are grateful to Y.Süzen for her kind support of this work.
+Author would like to express the gratitude to referees for constructive recommendations, making the computational work more robust and pointing out relevant literature.